If this 2.0 ppm dilution is still measurable, the detection limit for STD-B is less than 2 ppm. You cannot add more than 7 mL to a TT (it will overflow), so for a more dilute sample, you will have to begin with a smaller amount of STD-B. Adjust the pipette to withdraw just 0.10 mL of STD-B, and transfer this to a new TT. Calculate the volume of solvent solution that must be added to 0.10 mL of the 10.0 ppm STD-B drug standard to prepare a 1.0 ppm solution of STD-B, and enter your result below. Volume of solvent to produce a 1 ppm concentration mL Using an appropriate setting of the pipette, add this amount of solvent to the 0.10 mL of 10.0 ppm STD-B solution already in the TT. Then mix the solution with the TT mixer. Next
Answer #1

The question is about calculating the volume of solvent that must be added to 0.1 mL (V1) of 10 ppm (C1) solution to get final concentration of 1.0 ppm (C2)

Now,according to dilution formula

C1 * V1. = C2 * V2

Concentrated solution Dilute solution

V2 is final volume of diluted solution that will be formed

10 ppm * 0.1 mL. = 1 ppm * V2

V2 = (10 ppm * 0.1 mL)/ 1 ppm = 1 mL

So, 1 mL solution containing 0.1 mL of 10 ppm solution is of concentration 1 ppm.

Now, volume of final solution (V2) = volume of 10 ppm solution + volume of solvent

1 mL = 0.1 mL + volume of solvent

Volume of solvent = 1 mL - 0.1 mL = 0.9 mL

So, solvent to be added = 0.9 mL (Answer)